OpenStudy (anonymous):

Model the function rule y= -2x -3 with a table of values.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

@skullpatrol this is my first time taking Algebra 1

OpenStudy (skullpatrol):

Think of "x" as a symbol used to represent one or more numbers.

OpenStudy (anonymous):

@skullpatrol ok! But I still don't understand it

OpenStudy (skullpatrol):

Choose some values for that variable "x" and substitute them in for "x".

OpenStudy (skullpatrol):

Failure to follow the guidelines may lead to your suspension or banning from the site. If you see any violations, please report them by using a report abuse link. OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

y=-2x-3 (plug in random number) y=-2(3)-3 (plugged in 3 for x) y=-6-3 y=-9 (see what i did there?) Next, try plugging in another number for x, such as 4...etc,etc

OpenStudy (anonymous):

then you will start to see a pattern in the answers:)

OpenStudy (anonymous):

yes that help